OpenStudy (anonymous):

Question : 15 -Signals: \(y(t) = Real[x(t)]\) Check this system for Linearity and for Time-Invariance...

OpenStudy (dan815):

a system is saying to be time invariant if/ for input x(t) mapped to y(t) then x(t-k) is mapped to y(t-k) such a response means it is time invariant

OpenStudy (dan815):

Linearating L(c1x1+c2x2)=c1L(x1)+c2L(x2) it is called a linear transformation when the response of 2 inputs together is the sum of the respond of each input applied separetely

OpenStudy (dan815):

L(c1x)=L(x+x+...c1 times).... so the definition above also covers constant multiple case

OpenStudy (dan815):

2 or more* inputs*

OpenStudy (anonymous):

You are simply telling the definition or theoretical thing, but I have difficulty in solving it.. :)

OpenStudy (perl):

Linearating is not a word. maybe Linearizing

OpenStudy (dan815):

where i come from we call that linearting

OpenStudy (dan815):

@Miracrown

OpenStudy (anonymous):

Suppose \(\large x(t) = e^{j \omega t}\) a complex signal: Then : \[x(t) \rightarrow y(t) \\x(t) \rightarrow \Re[x(t)] = \cos( \omega t)\] Here, if I check homogeneity, then, let us say I multiply with \(j\), \[j \cdot x(t) \rightarrow \Re[j \cdot x(t)] = -\sin( \omega t) \ne j \cdot y(t)\]

OpenStudy (anonymous):

If I multiply by j in this case, the real part will become Imaginary and Imaginary can become Real, and that changes the output.. May be, this is the way to prove it Non-Linear.. :(

OpenStudy (anonymous):

@miracrown ..